ACCT B715  Case #3 ratios
Comparative Balance Sheets as of December 31 ($ in 000s)
2017 2016 2015
Assets
Cash $100 $450 $400
marketable securities 0 500 300
accounts receivable 1900 1600 1700
inventory 2200 1650 1600
current assets 4200 4200 4000
PPE 6500 3100 2800
Total Assets $10,700 7,300 6,800
Liabilities and Equity
Accounts payable $1,350 900 750
other payables 400 350 500
current liabilities 1750 1250 1250
long-term debt 3250 1100 1250
total liabilities 5000 2350 2500
Contributed capital 2200 2200 2200
retained earnings 3500 2750 2100
total equity 5700    4950 4300
Total liabilities and equity $10,700 7,300 6,800
Comparative Income Statements for Years ending Deccember 31 ($ in 000s)
2017 2016 2015
Net sales $12,500 $11,000 $10,500
Cost of goods sold
  Materials 3500 2400 1600
  Labor 2200 2700 3200
  Mf'g costs 2400 2200 2000
total CGS 8100 7300 6800
Gross Profit 4400 3700 3700
Selling expenses 1200 1050 1100
General & administrative exp 1200 1300 1200
Total operating expenses 2400 2350 2300
Net Income 2000 $1,350 $1,400
1. Compute the ratios we talked about in class--do the following for all three years:
current, quick, debt-to-equity, debt-to-assets, ROA and ROE
** net profit margin and gross profit margin as well
2. compute for 2017 and 2016: turnover and days for both A/R and Inventory.
3. Compare growth rate in receivable and inventory with the growth rate in sales.
Answer #1
Requirement 1
Sr. No. Ratio Formulas YEAR
2017 2016 2015
1 Current Ratio =Current asset/Current Liabilities =(4200/1750) =(4200/1250) =(4000/1250)
2.4 Times 3.36 Times 3.2 Times
2 Quick Ratio =Quick Asset/Quick Liabilities =(4200-2000)/1750 =(4200-1650)/1250 =(4000-1600)/1250
1.257143 Times 2.04 Times 1.92 Times
3 Debt to Equity =Total Liabilities/Total equity =5000/5700 =2350/4950 =2500/4300
0.877193 Times 0.474747 Times 0.581395 Times
4 Debt to Assets =Total Debt/Total assets =5000/10700 =2350/7300 =2500/6800
0.46729 Times 0.321918 Times 0.367647 Times
5 ROA =Net Income/Average Total assets =2000/(10700+7300)% =1350/(7300+6800)%
11.11111 % 9.574468 %
6 ROE =Net Income/Shareholder's equity =2000/5700% =1350/4950% =1400/4300%
35.08772 % 27.27273 % 32.55814 %
7 Net Profit Margin =Net Profit/Total Revenue =2000/12500% =1350/11000% =1400/10500%
16 % 12.27273 % 13.33333 %
8 Gross Profit Margin =Gross profit/ Total Revenue =4400/12500% =3700/11000% =3700/10500%
35.2 % 33.63636 % 35.2381 %
Requirement 2
1-a A/R Turnover Ratio =Net Credit Sales/Average A/R =12500/(1900+1600) =11000/(1600+1700)
3.571429 Times 3.333333 Times
1-b Days for A/R =(Ending A/R / Net Credit Sales) *365 =(1900/12500)*365 =(1600/11000)*365
55.48 Days 53.09091 Days
2-a Inventory Turnover Ratio =Cost of goods sold/Average Inventory =8100/(2200+1650) =(7300/(1650+1600)
2.103896 Times 2.246154 Times
2-b Days for Inventory =(Ending Inventory/Cost of goods sold)*365 =(2200/8100)*365 =(1650/7300)*365
99.1358 Days 82.5 Days
Requirement 3
1 Growth rate in sales =(12500-11000)/11000% =(11000-10500)/10500%
13.63636 % 4.761905 %
2 Growth rate in Receivable =(1900-1600)/1600% =(1600-1700)/1700%
18.75 % -5.88235 %
3 Growth rate in Inventory =(2200-1650)/1650% =(1650-1600)/1600%
33.33333 % 3.125 %